OpenStudy (anonymous):

verify the identity. 1-sinx/cosx = cosx/1+sinx

OpenStudy (anonymous):

you can "cross multiply" maybe and it should work out nicely

OpenStudy (anonymous):

you get \(\cos^2(x)=(1+\sin^2(x))(1-\sin^2(x))\) and since \[(1+\sin^2(x))(1-\sin^2(x))=1-\sin^2(x)\] you are pretty much done

OpenStudy (anonymous):

thanks :)